How can I calculate Ksp? All I have is KNO3 was dissolved in water and heated until the crystals dissolved. Then the solution was removed from heat and the crystals were allowed to reform- the temperature they reformed at was measured which was called "T of Dissolution". Then the test tube was heated again and the temperature at which the crystals reappeared was called "T of first crystals". These trials were repeated with varying mass of KNO3. Also- how to find molarity of KNO3?

Data: Trial 1

Mass KNO3: / T Dissolution / T First Crystals /Ksp?

2.499 g / 36.0 C / 31.2 C

2.499 / 35.1 C / 30.9 C

2.499 / 36.5 C / 31.0 C

Trial 2:

4.0 g / 63.3 C / 40.4 C

4.0 g / 62.6 C / 42.6 C

4.0 g / 60.4 C / 44.5 C

Trial 3

5.25 g / 60.0 C / 55.5 C

5.25 g / 63.7 C / 56.0 C

5.25 g / 62.3 C / 52.0 C

Trial 4

6.5 g / 63.0 C / 55.0 C

6.5 g / 62.4 C / 52.8 C

6.5 g / 63.6 C / 53.1 C

Trial 5

8.0 g / 80.6 C / 75.0 C

8.0 g / 79.9 C / 74.3 C

8.0 g / 80.3 C / 74.3
Answer #1

Solution:

When a salt dissolves in water it will dissociate into ions. In aqueous solution potassium nitrate (KNO3) dissociates according to the following reaction.

KNO3(s) ⇌ K + (aq) + NO3 - (aq)

Once a sufficient amount of K+ and NO3 - are in solution, the ions will recombine into solid potassium nitrate, KNO3. At this point a pair of ions will recombine for every pair that dissociates, and the concentrations of the ions in solution will remain constant. The reaction is said to be at equilibrium. The two-sided arrow indicates that the reaction is reversible and undissolved solid is in equilibrium with the dissolved ions. This is known as a saturated solution. The equilibrium constant expression, shown in equation (1), defines the equilibrium constant (Ksp) for KNO3 dissolving in water. Ksp = [K+ ][NO3 - ] The “sp” stands for the solubility product, and the brackets around the ions symbolize their molar concentration (Molarity, M or mol/L). The equilibrium constant is not only concerned with solubility but also is expressed as a product of the molar concentration of ions. The value for Ksp for a compound changes with the temperature. Ksp is different for different solutes. If ‘s’ is the solubility of of KNO3 then Ksp = [K+ ][NO3 - ] = s × s = s2

Ksp has no units. • Because the solubility of KNO3 changes when the temperature (T) changes, the Ksp is a function of the temperature and hence you are required to record the temperature of dissolution

Molarity is defined as

Molarity = moles of solute / liters of solution

Molar mass of KNO3 = 39.0983 + 14.0067 + 15.9994×3 = 101.1032 g/mol


mass of KNO3 taken / molar mass of KNO3 = Moles of KNO3

Molarity of KNO3 solution= moles of KNO3/ volume of the solution in litres

In general, Molarity (M) = (Weight of the solute/ Molar mass of the solute)× (1000/V) where V is the volume of the solution in mL